OpenStudy (anonymous):

Find all y such that the distance between the points (-6, -9) and (-4, y) is 28.

OpenStudy (anonymous):

Should you try Pythagorean's theorem?

OpenStudy (steve816):

Nope, use the distance formula.

OpenStudy (anonymous):

Like @steve816 said,use the distance formula :)

OpenStudy (anonymous):

Your distance formula is Pythagorean's...

OpenStudy (anonymous):

i did but i got sqrt y^2+18y+85=28. do i square both sides?

OpenStudy (anonymous):

@steve816 @RainbowBrony555

OpenStudy (anonymous):

A^2+B^2=C^2 where C is your distance Let a be x and b be y (-6- -4)^2+(-9 + y)^2= 28^2 You will have 2 answers for y. You could also graph it by hand...use a compass...radius=28